amy's age = a, her brother's age = b
Amy is 6 years younger than her brother -> a = b-6. In four years, she will be half as old as her brother -> a+4 = 1/2x(b+4).
How old is each child? a, b = ?

Answer by amalialaila(3)

a = b-6 ...(i)
a+4 = 1/2(b+4) ...(ii)
substitute equation (i) to (ii)
(b-6)+4 = 1/2(b+4)
b-2 = 1/2(b+4)
2(b-2) = b+4
2b-4 = b+4
unify the components with the same variable
2b-b = 4+4
b = 8, this is amy's brother age. substitute the number to equation (i)
a = 8-6
a = 2
amy's 2 years old and the brother is 8 years old.